# A Generative Approach to Control Complex Physical Systems

### v1

@@ 6 Conclusion
In this work, we have introduced DiffPhyCon, a novel methodology for controlling complex physical systems. It generates control sequences and state trajectories by jointly optimizing the generative energy and control objective. We further introduced prior reweighting to enable the discovery of control sequences that diverge significantly from training. Through comprehensive experiments, we demonstrated our method’s superior performance compared to classical, deep learning, and reinforcement learning baselines in challenging physical systems control tasks. We discuss future work in Appendix L and state social impact in M.
@@ A.3 Diffusion Models
Appendix B 1D Burgers’ Equation Visualization
@@
Figure 6: Visualizations results of 1D Burgers’ equation control under the FO-PC (full observation, partiall control) setting. The curve for the system state $\mathbf{u}_{t}$ of each time step $t=0,\cdots,10$ under control is plotted for our method (DiffPhyCon) and baselines. The $x$-axis is the spatial coordinate and the $y$-axis is the value of the system state.
From these visualization results in in Figure 6, Figure 7, and Figure 8, it can be observed that under the control of our DiffPhyCon, the system state could converge smoothly to the target state given different initial states, and the final state $\mathbf{u}_{T}$ always coincides with the target state. Furthermore, this observation is consistent under all three settings: FOPC, POFC, and POPC, implying that our DiffPhyCon is effective in addressing the partial observation and partial control challenges. In contrast, the baselines showed inferior results. Even the best baseline, BPPO, presents obvious mismatching with the target state on some samples.
Figure 7: Visualizations results of 1D Burgers’ equation control under the PO-FC (partiall observation, full control) setting. The curve for the system state $\mathbf{u}_{t}$ of each time step $t=0,\cdots,10$ under control is plotted for our method (DiffPhyCon) and baselines. The $x$-axis is the spatial coordinate and the $y$-axis is the value of the system state.
@@
Appendix C Additional Details for 1D Burgers’ Equation Control
C.1 Data Generation
@@
In the numerical simulation (using the ground-truth solver), a domain of $x=[0,1],~{}t=[0,1]$ is simulated. The space is discretized into $128$ grids and time into $10000$ steps. However, in the dataset, only 10 time stamps are stored. For the control sequence $\mathbf{w}$, its refreshing rate is $0.1^{-1}$, i.e., $\mathbf{w}(t,x),~{}t\in[0.1k,0.1(k+1)],~{}k\in\{0,..,9\}$ does not change with $t$. Therefore, the data size of each trajectory is $[11,128]$ for the state $\mathbf{u}$ and $[10,128]$ for the control $\mathbf{w}$.
@@
where each parameter is independently generated as follows: $b_{1,i}\sim U(0,1),~{}b_{2,i}\sim U(0,1),~{}\sigma_{1,i}\sim U(0.05,0.2),~{}\sigma_{2,i}\sim U(0.05,0.2)$, while $a_{1}\sim U(-1.5,1.5)$ and for $i\geq 2$, $a_{i}\sim U(-1.5,1.5)$ or 00 with equal probabilities. $\mathbf{u}(t,x),~{}(t\neq 0)$ is then numerically simulated (using the ground-truth solver) given $\mathbf{u}(0,x)$ and $\mathbf{w}(t,x)$ based on Eq. (16). The setting of the dataset generation is based on a previous work [23].
We generated $90000$ trajectories for the training set and $50$ for the testing set. Each trajectory takes up $32$KB space and the size of the dataset sums up to $2$GB.
C.2 Experimental Setting
During inference, alongside the control sequence $\mathbf{w}(t,x)$, our diffusion model generates states $\mu(t,x)$, and some models produce surrogate states $\mu(t,x)$ when feeding the control $\mathbf{w}(t,x)$ into the corresponding surrogate model. However, our reported evaluation metric $\mathcal{J}_{\text{actual}}$ is always computed by feeding the control $\mathbf{w}(t,x)$ into the ground truth numerical solver to get $\mathbf{u}_{\text{g.t.}}(t,x)$ and computed following Eq. (17).
Following are three different settings of our experiments.
C.2.1 Partial Observation, Full Control
@@
C.2.2 Full Observation, Partial Control
@@
C.2.3 Partial Observation, Partial Control
@@ C.3.1 DiffPhyCon-lite
In our experiments, we found that the control result is best when learning the conditional probability distribution of $p(\mathbf{w}_{[0,T-1]},\mathbf{u}_{[1,T-1]}\mid\mathbf{u}_{0},\mathbf{u}_{T})$ In summary, $\bm{\epsilon}_{\theta}$ takes in the current trajectory $\mathbf{u}$, control $\mathbf{w}$, step $k$, $\mathbf{u}_{0}$ and $\mathbf{u}_{T}$ as input, and predicts the noise of $\mathbf{u}$ and $\mathbf{w}$. Note that it is not trained to predict $\mathbf{u}_{0}$ and $\mathbf{u}_{T}$ which are used as a condition, but there are still model outputs at the corresponding locations for the data shape consistency across different design choices of DiffPhyCon-lite. The hyperparameters in different settings are listed in Table 4.
C.3.2 DiffPhyCon
@@ C.4 Training and Evaluation
During training, the $\mathbf{u}_{0}$ and $\mathbf{u}_{d}$ without noise are fed into the model and the model outputs at the corresponding locations are excluded from the loss. In the partial observation settings, the unobserved data is invisible to the model during both training and testing as introduced in Appendix C.2. We simply pad zero in the corresponding locations of the model input and also exclude these locations in the training loss. Therefore, the model only learns the correlation between the observed states and control sequences. In the partial control setting, we train our DiffPhyCon on the dataset with control being zero in $x\in[\frac{1}{4},\frac{3}{4}]$. In this way, the model naturally learns to output zero at the “non-controllable” locations.
@@
After generating the trajectory $\mathbf{u}$ and control sequence $\mathbf{w}$, we feed the control sequence $\mathbf{w}$ into the ground-truth solver and simulate the final state $\mathbf{u}$ given the generated $\mathbf{w}$ and the initial condition $\mathbf{u}_{0}$ directly drawn from the testing dataset. The solver is the same as the one used in data generation in Appendix C.1. Finally, we compute $\mathcal{J}_{\text{actual}}$ following Eq. (17). In the partial observation setting, the MSE is computed only on the observed region, and in the control setting, the generated control will first be set to zero in the uncontrolled region before being fed into the solver.
Appendix D Jellyfish Movement Dataset
@@
  • opening angels $\mathbf{w}$: shape $[\tilde{T}]$. For each step, we save the opening angle in radians.
@@
Appendix E Additional Details for 2D Jellyfish Movement Control
E.1 Dataset Preparation
Based on our generated dataset in Appendix D, we prepare training samples for the 2D jellyfish movement control task as follows. We use sliding time windows that contain $T=20$ successive time steps of states and boundaries as a sample, which corresponds to $T^{{}^{\prime}}=200$ original simulation steps and constitutes exactly a period of wing movement. In this way, each trajectory can produce 20 samples. Therefore, we get 6 million training samples in total. In each training sample, the initial and the final time steps share the same opening angle due to periodicity, which serves as the conditions for control. For each test trajectory, we select the opening angle of the jellyfish in the initial time and the initial states as the control condition for both the initial and final time and state initial condition.
Figure 9: Example of a flapping jellyfish in Lily-Pad simulator (a) and the corresponding curve of opening angle (b). The control signal is the opening angels $\mathbf{w}$ of the wings of a jellyfish.
E.2 Experimental Setting
E.2.1 Full Observation
@@ E.3.3 DiffPhyCon
E.4 Training, Inference, and Evaluation
@@
Inference. The pipeline of inference is shown in Figure 10. Both diffused variables $\mathbf{u}_{[0,T]}$ and $\mathbf{w}_{[0,T]}$ are initialized from Gaussian prior and gradually denoised from denoising step $k=1000$ to $k=0$ based on denoising networks and guidance. Because we introduce the boundary mask and offsets as auxiliary inputs, the model input and output are not consistent in shape. Thus we introduce a surrogate model (shown as "Boundary updater" block in Figure 10) to update boundary mask and offsets $b_{[0,T],k}$ for each denoising $k$. Specifically, at each time step $t\in[0,T]$, $b_{t,k}$ is estimated by the initial boundary mask and offsets $b_{0}$, and the difference of opening angle $\mathbf{w}_{t,k}-\mathbf{w}_{0}$ from time step 00 to $t$, which is presented in the right part (after "Denoise under guidance") of Figure 10. Details about this surrogate model are presented in H.3. Notice that although this surrogate model is trained on noise-free data, we do not worry too much about its generalization to the noisy scalar $\mathbf{w}_{t,k}$ in inference because the estimated $\mathbf{w}_{t,k}$ does not deviate from the normalized range of noisy free $\mathbf{w}_{t}$ too much.
@@
As for guidance, we use a surrogate force model to approximate the force of fluid on jellyfish. This model is detailed in Subsection H.2. In denoising step $k$, its input consists of two parts: the first one is the noise-free state $\hat{\mathbf{u}}_{[0,T]}$ estimated from $\mathbf{u}_{[0,T],k}$ by Eq. (6); the second one is the noise-free boundary mask and offsets $\hat{b}_{[0,T],k}$ estimated from noise-free $\hat{\mathbf{w}}_{[0,T]}$ by the surrogate model to update boundaries, where $\hat{\mathbf{w}}_{[0,T]}$ is also estimated from $\mathbf{w}_{[0,T],k}$ as in Eq. (6). The model output is force. Here we only use the horizontal force. Notice that the force could be computed via the surrogate force model no matter whether states are fully or partial observation in that force is irrelevant to the velocity of the fluid. The control objective $\mathcal{J}$ in Eq. (19) is computed as a summation of force and $R(\hat{\mathbf{w}}_{[0,T]})$. We fix $\zeta=1000$ as a default setting in Eq. (19) because this value can achieve a balance between scales of the average speed and the regularizer $R(\mathbf{w})$. We also study the Pareto performance of varying $\zeta$ in Table 3. Then the gradients of the objective $\mathcal{J}$ in terms of $\hat{\mathbf{u}}_{[0,T]}$ and $\hat{\mathbf{w}}_{[0,T]}$ are computed and used in guidance. For DiffPhyCon-lite, these gradients are substracted from $[\mathbf{u}_{[0,T],k},\mathbf{w}_{[0,T],k}]$ to generate $[\mathbf{u}_{[0,T],k-1},\mathbf{w}_{[0,T],k-1}]$. For DiffPhyCon, an additional term of noise $(\gamma-1)\bm{\epsilon}_{\phi}$ predicted from the denoising network of the marginal distribution $p(\mathbf{w}|\mathbf{u}_{0},\mathbf{w}_{0})$ should also be subtracted, as shown in the upper left part of Figure 10. The effect of this term is controlled by the scale of the hyperparameter $\gamma$, as studied in Appendix J. The inference is performed on one NVIDIA Tesla-V100 GPU with 32 GB memory for about 3 hours for 50 testing samples.

### v3

@@ 5 Conclusion
In this work, we have introduced DiffPhyCon, a novel methodology for controlling complex physical systems. It generates control sequences and state trajectories by jointly optimizing the generative energy and control objective. We further introduced prior reweighting to enable the discovery of control sequences that diverge significantly from training. Through comprehensive experiments, we demonstrated our method’s superior performance compared to classical, deep learning, and reinforcement learning baselines in challenging physical systems control tasks. We discuss limitation and future work in Appendix N and state social impact in O.
6 Acknowledgment
We thank Yuchen Yang for insightful discussions on theoretical analysis. We thank the anonymous reviewers for providing valuable feedback on our manuscript. We also gratefully acknowledge the support of Westlake University Research Center for Industries of the Future and Westlake University Center for High-performance Computing. The content is solely the responsibility of the authors and does not necessarily represent the official views of the funding entities.
Appendix
[…]
@@ A.3 Diffusion Models
Appendix B Theoretical Analysis of Prior Reweighting
Proof of Theorem 3.1.
$E(\gamma)=$ $\int\mathbb{I}_{Q(\varepsilon)}(\mathbf{u},\mathbf{w})p_{\gamma}(u,w|Y=1)\mathrm{d}(\mathbf{u},\mathbf{w})$
$=$ $\int\mathbb{I}_{Q(\varepsilon)}(\mathbf{u},\mathbf{w})\frac{p(Y=1|\mathbf{u},\mathbf{w})p_{\gamma}(\mathbf{u},\mathbf{w})}{p(Y=1)}\mathrm{d}(\mathbf{u},\mathbf{w})$
$=$ $\frac{\mathbb{E}_{(\mathbf{u},\mathbf{w})}[\mathbb{I}_{Q(\varepsilon)}(\mathbf{u},\mathbf{w})p(Y=1|\mathbf{u},\mathbf{w})]}{\mathbb{E}_{(\mathbf{u},\mathbf{w})}[p(Y=1|\mathbf{u},\mathbf{w})]}$
$=$ $\frac{\mathbb{E}_{(\mathbf{u},\mathbf{w})\}[\mathbb{I}_{Q(\varepsilon)}(\mathbf{u},\mathbf{w})p(Y=1|\mathbf{u},\mathbf{w})]}}{\mathbb{E}_{\mathbf{u},\mathbf{w}}[\mathbb{I}_{Q(\varepsilon)}(\mathbf{u},\mathbf{w})p(Y=1|\mathbf{u},\mathbf{w})]+\mathbb{E}_{\mathbf{u},\mathbf{w}}[\mathbb{I}_{Q(\varepsilon)^{c}}(\mathbf{u},\mathbf{w})p(Y=1|\mathbf{u},\mathbf{w})]}$
$=$ $\frac{1}{1+\frac{\mathbb{E}_{\mathbf{u},\mathbf{w}}[\mathbb{I}_{Q(\varepsilon)^{c}}(\mathbf{u},\mathbf{w})p(Y=1|\mathbf{u},\mathbf{w})]}{\mathbb{E}_{\mathbf{u},\mathbf{w}}[\mathbb{I}_{Q(\varepsilon)}(\mathbf{u},\mathbf{w})p(Y=1|\mathbf{u},\mathbf{w})]}}$
Define
$G(\gamma)=$ $\frac{\mathbb{E}_{\mathbf{u},\mathbf{w}}[\mathbb{I}_{Q(\varepsilon)}(\mathbf{u},\mathbf{w})p(Y=1|\mathbf{u},\mathbf{w})]}{\mathbb{E}_{\mathbf{u},\mathbf{w}}[\mathbb{I}_{Q(\varepsilon)^{c}}(\mathbf{u},\mathbf{w})p(Y=1|\mathbf{u},\mathbf{w})]}$
$=$ $\frac{\mathbb{E}_{w}[\mathbb{E}_{u}[\mathbb{I}_{Q(\varepsilon)}(\mathbf{u},\mathbf{w})p(Y=1|\mathbf{u},\mathbf{w})|\mathbf{w}]]}{\mathbb{E}_{w}[\mathbb{E}_{u}[\mathbb{I}_{Q(\varepsilon)^{c}}(\mathbf{u},\mathbf{w})p(Y=1|\mathbf{u},\mathbf{w})|\mathbf{w}]]}$
$=$ $\frac{\int\mathbb{E}_{u}[\mathbb{I}_{Q(\varepsilon)}(\mathbf{u},\mathbf{w})p(Y=1|\mathbf{u},\mathbf{w})|\mathbf{w}]p^{\gamma}(\mathbf{w})\mathrm{d}\mathbf{w}}{\int\mathbb{E}_{u}[\mathbb{I}_{Q(\varepsilon)^{c}}(\mathbf{u},\mathbf{w})p(Y=1|\mathbf{u},\mathbf{w})|\mathbf{w}]p^{\gamma}(\mathbf{w})\mathrm{d}\mathbf{w}}$
Then
$E(\gamma)=\frac{1}{1+\frac{1}{G(\gamma)}}$. Since $G(\gamma)>0$, $E(\gamma)$ and $G(\gamma)$ have the same monotonicity.
$G^{\prime}(\gamma)=$ $\frac{\int\mathbb{E}_{u}[\mathbb{I}_{Q(\varepsilon)}(\mathbf{u},\mathbf{w})p(Y=1|\mathbf{u},\mathbf{w})|\mathbf{w}]p^{\gamma}(\mathbf{w})\ln(p(\mathbf{w}))\mathrm{d}\mathbf{w}\mathbb{E}_{\mathbf{u},\mathbf{w}}[\mathbb{I}_{Q(\varepsilon)^{c}}(\mathbf{u},\mathbf{w})p(Y=1|\mathbf{u},\mathbf{w})]}{(\mathbb{E}_{\mathbf{u},\mathbf{w}}[\mathbb{I}_{Q(\varepsilon)^{c}}(\mathbf{u},\mathbf{w})p(Y=1|\mathbf{u},\mathbf{w})])^{2}}$
$-\frac{\mathbb{E}_{\mathbf{u},\mathbf{w}}[\mathbb{I}_{Q(\varepsilon)}(\mathbf{u},\mathbf{w})p(Y=1|\mathbf{u},\mathbf{w})]\int\mathbb{E}_{u}[\mathbb{I}_{Q(\varepsilon)^{c}}(\mathbf{u},\mathbf{w})p(Y=1|\mathbf{u},\mathbf{w})|\mathbf{w}]p^{\gamma}(\mathbf{w})\ln(p(\mathbf{w}))\mathrm{d}\mathbf{w}}{(\mathbb{E}_{\mathbf{u},\mathbf{w}}[\mathbb{I}_{Q(\varepsilon)^{c}}(\mathbf{u},\mathbf{w})p(Y=1|\mathbf{u},\mathbf{w})])^{2}}$
$=$ $\frac{\mathbb{E}_{\mathbf{u},\mathbf{w}}[\mathbb{I}_{Q(\varepsilon)}(\mathbf{u},\mathbf{w})p(Y=1|\mathbf{u},\mathbf{w})\ln(p(\mathbf{w}))]\mathbb{E}_{\mathbf{u},\mathbf{w}}[\mathbb{I}_{Q(\varepsilon)^{c}}(\mathbf{u},\mathbf{w})p(Y=1|\mathbf{u},\mathbf{w})]}{(\mathbb{E}_{\mathbf{u},\mathbf{w}}[\mathbb{I}_{Q(\varepsilon)^{c}}(\mathbf{u},\mathbf{w})p(Y=1|\mathbf{u},\mathbf{w})])^{2}}$
$-\frac{\mathbb{E}_{\mathbf{u},\mathbf{w}}[\mathbb{I}_{Q(\varepsilon)}(\mathbf{u},\mathbf{w})p(Y=1|\mathbf{u},\mathbf{w})]\mathbb{E}_{\mathbf{u},\mathbf{w}}[\mathbb{I}_{Q(\varepsilon)^{c}}(\mathbf{u},\mathbf{w})p(Y=1|\mathbf{u},\mathbf{w})\ln(p(\mathbf{w}))]}{(\mathbb{E}_{\mathbf{u},\mathbf{w}}[\mathbb{I}_{Q(\varepsilon)^{c}}(\mathbf{u},\mathbf{w})p(Y=1|\mathbf{u},\mathbf{w})])^{2}}$
By definition, $F(\gamma)$ is a positive multiple of $G^{\prime}(\gamma)$, which implies our conclusion:
  • If $F(1)<0$, then $G^{\prime}(1)<0$ and thus $E(\gamma)$ decreases around 1. Hence, there exists $\gamma_{-}<1$, s.t., $E(\gamma_{-})>E(1)$, thus (i) holds;
  • otherwise, for similar reason, (ii) holds.
Remark: Here $F(\gamma)$ can be interpreted as some kind of difference between "entropies" in $Q(\varepsilon)^{c}$ and $Q(\varepsilon)$. When $F(1)<0$, it means that $Q(\varepsilon)^{c}$ has higher "entropies", implying that the training trajectories are far from optimal. As a result, we may need to flatten the distribution of training trajectories, which corresponds to using the prior reweighting technique with $\gamma<1$. Since this is the most common case, we usually set $\gamma<1$.
Appendix C Visualization Results
C.1 1D Burgers’ Equation Visualization
@@ Appendix C Visualization Results
Figure 7: Visualizations results of 1D Burgers’ equation control under the FO-PC (full observation, partial control) setting. The curve for the system state $\mathbf{u}_{t}$ of each time step $t=0,\cdots,10$ under control is plotted for our method (DiffPhyCon) and baselines. The $x$-axis is the spatial coordinate and the $y$-axis is the value of the system state.
From these visualization results in Figure 7, Figure 8, and Figure 9, it can be observed that under the control of our DiffPhyCon, the system state could converge smoothly to the target state given different initial states, and the final state $\mathbf{u}_{T}$ always coincides with the target state. Furthermore, this observation is consistent under all three settings: FOPC, POFC, and POPC, implying that our DiffPhyCon is effective in addressing the partial observation and partial control challenges. In contrast, the baselines showed inferior results. Even the best baseline, BPPO, presents obvious mismatching with the target state on some samples.
Figure 8: Visualizations results of 1D Burgers’ equation control under the PO-FC (partial observation, full control) setting. The curve for the system state $\mathbf{u}_{t}$ of each time step $t=0,\cdots,10$ under control is plotted for our method (DiffPhyCon) and baselines. The $x$-axis is the spatial coordinate and the $y$-axis is the value of the system state.
@@
C.2 2D jellyfish control Visualization
We present more simulation results of our method in Figure 10. Each line represents an example from the test dataset. We plot five snapshots of boundary and fluid field for each example.
Figure 10: More examples of 2D jellyfish simulation controlled by our method.
C.3 2D Smoke Control Visualization
We present fluid states and control signals generated by our method in Figure 11. Each line represents an example from the test dataset. We plot six snapshots for each example.
Figure 11: Examples of 2D smoke control results by our method. We present three randomly selected test examples. For each example, we show the generated smoke density map and control force fields in horizontal and vertical directions. Each row depicts six frames of movement. The smoke density in the first row corresponds to that in Figure 6.
Appendix D Additional Details for 1D Burgers’ Equation Control
D.1 Data Generation
@@
In the numerical simulation (using the ground-truth solver), a domain of $x\in[0,1],~{}t\in[0,1]$ is simulated. The space is discretized into $128$ grids and time into $10000$ steps. However, in the dataset, only 10 time stamps are stored. For the control sequence $\mathbf{w}$, its refreshing rate is $0.1^{-1}$, i.e., $\mathbf{w}(t,x),~{}t\in[0.1k,0.1(k+1)],~{}k\in\{0,..,9\}$ does not change with $t$. Therefore, the data size of each trajectory is $[11,128]$ for the state $\mathbf{u}$ and $[10,128]$ for the control $\mathbf{w}$.
@@
where each parameter is independently generated as follows: $b_{1,i}\sim U(0,1),~{}b_{2,i}\sim U(0,1),~{}\sigma_{1,i}\sim U(0.05,0.2),~{}\sigma_{2,i}\sim U(0.05,0.2)$, while $a_{1}\sim U(-1.5,1.5)$ and for $i\geq 2$, $a_{i}\sim U(-1.5,1.5)$ or 00 with equal probabilities. $\mathbf{u}(t,x),~{}(t\neq 0)$ is then numerically simulated (using the ground-truth solver) given $\mathbf{u}(0,x)$ and $\mathbf{w}(t,x)$ based on Eq. (16). The setting of the dataset generation is based on a previous work [24]. We generated $90000$ trajectories for the training set and $50$ for the testing set. Each trajectory takes up $32$KB space and the size of the dataset sums up to $2$GB.
D.2 Experimental Setting
During inference, alongside the control sequence $\mathbf{w}(t,x)$, our diffusion model generates states $\mu(t,x)$, and some models produce surrogate states $\mu(t,x)$ when feeding the control $\mathbf{w}(t,x)$ into the corresponding surrogate model. However, our reported evaluation metric $\mathcal{J}_{\text{actual}}$ is always computed by feeding the control $\mathbf{w}(t,x)$ into the ground truth numerical solver to get $\mathbf{u}_{\text{g.t.}}(t,x)$ and computed following Eq. (17). Followings are three different settings of our experiments.
D.2.1 Partial Observation, Full Control
@@
D.2.2 Full Observation, Partial Control
@@
D.2.3 Partial Observation, Partial Control
@@ D.3.1 DiffPhyCon-lite
In our experiments, we found that the control result is best when learning the conditional probability distribution of $p(\mathbf{w}_{[0,T-1]},\mathbf{u}_{[1,T-1]}\mid\mathbf{u}_{0},\mathbf{u}_{T})$. In summary, $\bm{\epsilon}_{\theta}$ takes in the current trajectory $\mathbf{u}$, control $\mathbf{w}$, step $k$, $\mathbf{u}_{0}$ and $\mathbf{u}_{T}$ as input, and predicts the noise of $\mathbf{u}$ and $\mathbf{w}$. Note that it is not trained to predict $\mathbf{u}_{0}$ and $\mathbf{u}_{T}$ which are used as a condition, but there are still model outputs at the corresponding locations for the data shape consistency across different design choices of DiffPhyCon-lite. The hyperparameters in different settings are listed in Table 5.
D.3.2 DiffPhyCon
@@ D.4 Training and Evaluation
During training, the $\mathbf{u}_{0}$ and $\mathbf{u}_{d}$ without noise are fed into the model and the model outputs at the corresponding locations are excluded from the loss. In the partial observation settings, the unobserved data is invisible to the model during both training and testing as introduced in Appendix D.2. We simply pad zero in the corresponding locations of the model input and also exclude these locations in the training loss. Therefore, the model only learns the correlation between the observed states and control sequences. In the partial control setting, we train our DiffPhyCon on the dataset with control being zero in $x\in[\frac{1}{4},\frac{3}{4}]$. In this way, the model naturally learns to output zero at the “non-controllable” locations.
@@
After generating the trajectory $\mathbf{u}$ and control sequence $\mathbf{w}$, we feed the control sequence $\mathbf{w}$ into the ground-truth solver and simulate the final state $\mathbf{u}$ given the generated $\mathbf{w}$ and the initial condition $\mathbf{u}_{0}$ directly drawn from the testing dataset. The solver is the same as the one used in data generation in Appendix D.1. Finally, we compute $\mathcal{J}_{\text{actual}}$ following Eq. (17). In the partial observation setting, the MSE is computed only on the observed region, and in the control setting, the generated control will first be set to zero in the uncontrolled region before being fed into the solver.
Appendix E Jellyfish Movement Dataset
@@
  • opening angles $\mathbf{w}$: shape $[\tilde{T}]$. For each step, we save the opening angle in radians.
@@
Appendix F Additional Details for 2D Jellyfish Movement Control
F.1 Dataset Preparation
Based on our generated dataset in Appendix E, we prepare training samples for the 2D jellyfish movement control task as follows. We use sliding time windows that contain $T=20$ successive time steps of states and boundaries as a sample, which corresponds to $T^{{}^{\prime}}=200$ original simulation steps and constitutes exactly a period of wing movement. In this way, each trajectory can produce 20 samples. Therefore, we get 6 million training samples in total. In each training sample, the initial and the final time steps share the same opening angle due to periodicity, which serves as the conditions for control. For each test trajectory, we select the opening angle of the jellyfish in the initial time and the initial states as the control condition for both the initial and final time and state initial condition.
Figure 12: Example of a flapping jellyfish in Lily-Pad simulator (a) and the corresponding curve of opening angle (b). The control signal is the opening angles $\mathbf{w}$ of the wings of a jellyfish.
F.2 Experimental Setting
F.2.1 Full Observation
@@ F.3.3 DiffPhyCon
F.4 Training, Inference, and Evaluation
@@
Inference. The pipeline of inference is shown in Figure 13. Both diffused variables $\mathbf{u}_{[0,T]}$ and $\mathbf{w}_{[0,T]}$ are initialized from Gaussian prior and gradually denoised from denoising step $k=1000$ to $k=0$ based on denoising networks and guidance. Because we introduce the boundary mask and offsets as auxiliary inputs, the model input and output are not consistent in shape. Thus we introduce a surrogate model (shown as "Boundary updater" block in Figure 13) to update boundary mask and offsets $b_{[0,T],k}$ for each denoising $k$. Specifically, at each time step $t\in[0,T]$, $b_{t,k}$ is estimated by the initial boundary mask and offsets $b_{0}$, and the difference of opening angle $\mathbf{w}_{t,k}-\mathbf{w}_{0}$ from time step 00 to $t$, which is presented in the right part (after "Denoise under guidance") of Figure 13. Details about this surrogate model are presented in J.3. Notice that although this surrogate model is trained on noise-free data, we do not worry too much about its generalization to the noisy scalar $\mathbf{w}_{t,k}$ in inference because the estimated $\mathbf{w}_{t,k}$ does not deviate from the normalized range of noisy free $\mathbf{w}_{t}$ too much.
@@
As for guidance, we use a surrogate force model to approximate the force of fluid on jellyfish. This model is detailed in Subsection J.2. In denoising step $k$, its input consists of two parts: the first one is the noise-free state $\hat{\mathbf{u}}_{[0,T]}$ estimated from $\mathbf{u}_{[0,T],k}$ by Eq. (6); the second one is the noise-free boundary mask and offsets $\hat{b}_{[0,T],k}$ estimated from noise-free $\hat{\mathbf{w}}_{[0,T]}$ by the surrogate model to update boundaries, where $\hat{\mathbf{w}}_{[0,T]}$ is also estimated from $\mathbf{w}_{[0,T],k}$ as in Eq. (6). The model output is force. Here we only use the horizontal force. Notice that the force could be computed via the surrogate force model no matter whether states are fully or partial observation in that force is irrelevant to the velocity of the fluid. The control objective $\mathcal{J}$ in Eq. (19) is computed as a summation of force and $R(\hat{\mathbf{w}}_{[0,T]})$. We fix $\zeta=1000$ as a default setting in Eq. (19) because this value can achieve a balance between scales of the average speed and the regularizer $R(\mathbf{w})$. We also study the Pareto performance of varying $\zeta$ in Table 3. Then the gradients of the objective $\mathcal{J}$ in terms of $\hat{\mathbf{u}}_{[0,T]}$ and $\hat{\mathbf{w}}_{[0,T]}$ are computed and used in guidance. For DiffPhyCon-lite, these gradients are substracted from $[\mathbf{u}_{[0,T],k},\mathbf{w}_{[0,T],k}]$ to generate $[\mathbf{u}_{[0,T],k-1},\mathbf{w}_{[0,T],k-1}]$. For DiffPhyCon, an additional term of noise $(\gamma-1)\bm{\epsilon}_{\phi}$ predicted from the denoising network of the marginal distribution $p(\mathbf{w}|\mathbf{u}_{0},\mathbf{w}_{0})$ should also be subtracted, as shown in the upper left part of Figure 13. The effect of this term is controlled by the scale of the hyperparameter $\gamma$, as studied in Appendix L. The inference is performed on one NVIDIA Tesla-V100 GPU with 32 GB memory for about 3 hours for 50 testing samples.